# A multiple imputation approach to distinguish curative from life-prolonging effects in the presence of missing covariates

### v1

@@ 5 Case study
Previous study (Musta et al., 2022) showed that the histologic response is strongly prognostic for the cure fraction, while for the uncured patients, there is no clear indication that good histologic response is associated with longer PFS. On the other hand, there is no evidence of a statistically significant effect of the allocated regimen on the cure fraction or the survival rates, suggesting that the intensified treatment might not be associated with better cure chances and longer PFS.
We fit a Cox proportional hazards cure model to the complete data using the variables histological response and sex as predictors in the incidence part and histological response and assigned treatment as predictors in the latency part of the mixture cure model. Details regarding the test for the sufficient follow-up assumption are provided in Appendix C. We performed multiple imputation of the histological response using both the exact and the approximated approaches, as this variable suffers from $\sim 12$% missing values. Figure 2 illustrates the results of the Cox proportional hazards cure model for time-to-osteosarcoma recurrence.
@@
The only minor difference is that confidence intervals tend to be narrower for the imputation approaches than for complete-case analysis. The most notable difference between the imputation and complete-case fits is in the estimates of the model intercept, mainly in the width of his confidence interval, which is much wider in the complete-case setting.
@@ 6 Discussion
Our work on imputation methods for mixture cure models in the presence of missing covariate values emphasizes the importance of flexible imputation techniques. In this paper, we focus on improving the methodology for imputing partially observed covariates by accommodating potentially distinct sets of covariates for the cure probability and the survival of uncured patients.
Results show that both exact and approximate imputation approaches produce similar results. Notably, the imputation methods yield narrower confidence intervals compared to the complete-case analysis, enhancing the precision of our estimates.
@@
The simulation results suggest that the proposed methods are robust to the presence of missing values, particularly under the Missing at Random (MAR) scenario, where they outperform the complete-case analysis. This highlights the effectiveness of multiple imputation in handling complex missing data patterns, preserving empirical coverage, and reducing Mean Squared Error (MSE). The exact imputation method slightly outperforms the approximate method in terms of confidence interval width, yielding more precise estimates. However, when the model is misspecified, i.e. the covariate with missing values is wrongly included in both latency and incidence components, the approximate method has better coverage probabilities. These findings also point to the advantage of differentiating covariates between incidence and latency, providing a more nuanced and accurate analysis.
@@
For the BO06 clinical trial data on osteosarcoma, the application of the Cox proportional hazards cure model reveals several key findings. First, histological response emerges as a strong prognostic factor for the cure fraction, indicating that patients who respond well to initial treatment are more likely to be cured. However, again there is no clear indication that a good histological response is associated with longer progression-free survival (PFS) for uncured patients. This underscores the importance of having the possibility of differentiating between factors that influence cure probability from those that affect survival among uncured patients.
Overall, the enhanced methodology for multiple imputation, accommodating different sets of covariates for incidence and latency, proves to be a valuable tool in the analysis of survival data with missing covariates. This approach not only improves the accuracy and reliability of the estimates but also provides a more nuanced understanding of the factors influencing both cure probability and survival of uncured patients in osteosarcoma. The flexibility in covariate selection allows for more tailored and precise modelling, ultimately enhancing the interpretability and applicability of the results in clinical practice.
@@
As a further development of this work, one might consider an alternative way of performing multiple imputation based on the observed likelihood (instead of the complete one), which avoids considering the latent cure status as an additional variable with missing values. When there is only one missing variable, this would avoid the chained equations since there would be no need to impute G. However, such an approach would require the estimation of the parameters $\alpha$ and $\beta$ in Step 2 of Algorithm 1 by fitting a mixture cure model and the variance needs to be estimated via computationally intensive bootstrap. Hence, we did not explore this method further in this work.
@@ B.2 Incomplete covariate only in the latency
We test the sufficient follow-up assumption for the osteosarcoma data analyzed in Section 5. Currently there exist no test uniformly over all covariate values so we apply the tests proposed by Xie et al. [Xie et al., 2023] and Yuen and Musta [Yuen and Musta, 2024] separately for patients with good and poor histologic response, since that is the main prognostic factor for the cure probability.
The test proposed by Xie et al. [Xie et al., 2023] considers the null hypothesis of sufficient follow-up $H_{0}:\tau_{F_{u}}\leq\tau_{C}$, where $\tau_{F_{u}}$ and $\tau_{C}$ denote, respectively, the right end points of the distribution of the event times for the uncured patients and of the censoring times. For patients with good histologic response, such test does not reject the null hypothesis of sufficient follow-up since $2(Y_{(n)}-\tilde{Y}_{(m)})<Y_{(n)}$, where $Y_{(n)}$ denotes the largest observed time and $\tilde{Y}_{(m)}$ denotes the largest observed event time. For patients with poor histologic response, the p-value of the test is 0.038 meaning that the null hypothesis is rejected at level 0.05.
However, it is argued by Yuen and Musta [Yuen and Musta, 2024] that one could instead test a more relaxed notion of practically sufficient follow-up. The null hypothesis is that the follow-up is insufficient $\tilde{H}_{0}:\tau_{C}\leq q_{1-\epsilon}$, where $q_{1-\epsilon}$ denotes the $(1-\epsilon)$-quantile of $F_{u}$ and $\epsilon$ is a small number chosen by the user. It essentially tests whether there is less that $\epsilon$ probability for the event to happen after the end of the study. In practice this would be considered as a sufficiently long follow-up if $\epsilon$ is around $1$-$2\%$. The test requires also the choice of a parameter $\tau$ such that the chances of the event happening after $\tau$ are negligible, i.e. a study of length $\tau$ would have been certainly sufficient. We apply the test for two different choices of $\epsilon$ and $\tau$ and p-values are given in Table 4. The results indicate that the follow-up is clearly sufficient for patients with good histologic response and borderline sufficient for those with poor histologic response.

### v2

@@ 5 Case study
Our purpose is to just perform a sensitivity analysis of the results with respect to the method used in handling the missing data. We fit a Cox proportional hazards cure model to the complete data, selecting variables for the incidence and latency components using a combination of clinical knowledge from previous studies (Musta et al., 2022) and statistical considerations. We included histological response and assigned treatment as predictors in the latency component of the model. Conditioned on this choice, we selected sex and histological response for the incidence part of the model, as these variables were strongly associated with the probability of being cured. Notably, these variables would also have been selected using the penalized likelihood approach proposed by Masud et al. (2018). Details regarding the test for the sufficient follow-up assumption are provided in Appendix C. As histological response suffers from $\sim 12$% missing values, we performed multiple imputation of this variable using both the exact and the approximate approaches. Figure 2 illustrates the results of the Cox proportional hazards cure model for time-to-osteosarcoma recurrence.
@@
The only minor difference is that confidence intervals tend to be narrower for the imputation approaches than for complete-case analysis. The most notable difference between the imputation and complete-case fits is in the estimates of the model intercept, mainly in the width of his confidence interval, which is much wider in the complete-case setting. This behaviour is consistent with the results observed in the simulations.
As for the simulation in Scenario C, we also compare our extended method with the approach proposed by Beesley et al., incorporating an additional variable derived from the interaction term of histologic response and treatment.
Table 4 shows the results from the comparison between a full model, which includes all covariates in both the incidence and latency components, and a reduced model, obtained by selecting variables in the two parts.
Table 4: Comparison of Cox PH cure model estimates with a full and a reduced model for the analysis of osteosarcoma recurrence.
Full model Reduced model
Complete-case OR1 95% CI1 HR1 95% CI1 OR1 95% CI1 HR1 95% CI1
Intercept 3.24 1.90, 5.50 2.56 1.65, 3.99
Histological response
   Poor
   Good 0.19 0.09, 0.42 0.81 0.44, 1.51 0.29 0.17, 0.52 0.87 0.56, 1.35
Sex
   Female
   Male 1.48 1.01, 2.18 1.16 0.87, 1.54 1.58 1.09, 2.28
Allocated treatment
   Reg-C
   Reg-DI 0.53 0.25, 1.10 0.90 0.57, 1.42 0.81 0.55, 1.20
Treatment*Hist. resp.
   Good.Reg-DI 2.67 0.92, 7.74 0.99 0.42, 2.31
Exact
Intercept 2.45 1.37, 4.41 1.85 1.14, 3.00
Histological response
   Poor
   Good 0.19 0.09, 0.42 0.81 0.44, 1.52 0.29 0.17, 0.52 0.87 0.56, 1.35
Sex
   Female
   Male 1.74 1.01, 3.00 1.24 0.84, 1.82 1.91 1.14, 2.22
Allocated treatment
   Reg-C
   Reg-DI 0.53 0.25, 1.10 0.90 0.57, 1.42 0.82 0.55, 1.14
Treatment*Hist. resp.
   Good.Reg-DI 2.67 0.92, 7.75 0.99 0.42, 2.31
Approximate
Intercept 2.12 1.24, 3.61 1.84 1.19, 2.86
Histological response
   Poor
   Good 0.24 0.11, 0.49 0.90 0.50, 1.62 0.34 0.20, 0.57 0.93 0.62, 1.39
Sex
   Female
   Male 1.68 1.01, 2.80 1.18 0.82, 1.70 1.79 1.11, 2.37
Allocated treatment
   Reg-C
   Reg-DI 0.75 0.37, 1.53 0.82 0.53, 1.26 0.80 0.57, 1.12
Treatment*Hist. resp.
   Good.Reg-DI 2.12 0.75, 6.03 0.95 0.43, 2.11
1 OR: Odds Ratio, CI: Confidence Interval, HR: Hazard Ratio
The estimates from the saturated model show a wider uncertainty in parameter estimates. This is evident in the broader confidence intervals, especially for the intercept, sex, and treatment effects, indicating that the inclusion of unnecessary covariates introduces additional variability and reduces the precision of the estimates. In contrast, for the non-saturated models, both Exact and Approximate, the confidence intervals are narrower. The odds ratios for "Good" histological response are consistent across both imputation methods, but the confidence intervals are tighter in the Exact method, as observed in the simulation study. Overall, we observe that the precision of the parameter estimates improves when we limit the covariates to those that are meaningful in the context of the disease and treatment under study. This suggests that distinguishing between the covariates affecting the probability of being cured and those influencing the survival of uncured patients may be crucial for more efficient modelling.
@@ 6 Discussion
Our work on imputation methods for mixture cure models in the presence of missing covariate values emphasises the importance of flexible imputation techniques. In this paper, we investigate the methodology for imputing partially observed covariates in a mixture cure model by accommodating potentially distinct sets of covariates for the cure probability and the survival of uncured patients. This extension is motivated by both theoretical considerations - such as the identifiability benefits of distinct covariate sets - and practical challenges that arise when including all available covariates to both submodels. Importantly, we release the R-code to implement the proposed approach.
@@
The results show that both exact and approximate imputation approaches produce similar results when the model is correctly specified. In particular, imputation methods yield narrower confidence intervals compared to the complete case analysis, enhancing the precision of our estimates. The simulation results suggest that the proposed methods are robust to the presence of missing values, particularly under the MAR scenario, where they outperform the complete case analysis. This highlights the effectiveness of multiple imputation in handling complex missing data patterns, preserving empirical coverage, and reducing MSE. The exact imputation method slightly outperforms the approximate method in terms of confidence interval width, yielding more precise estimates. However, when the model is misspecified, i.e. certain covariates are wrongly included in the incidence or latency components, the approximate method has higher coverage probabilities, hence it is a safer and more robust choice for practical applications. Furthermore, the results illustrate the advantage of differentiating covariates between incidence and latency, since wrongly including all covariates in both components leads to wider confidence intervals and worse performance of the exact approach.
@@
For data from the BO06 clinical trial on osteosarcoma, the application of the Cox proportional hazards cure model reveals several key findings. First, as observed in the complete case analysis, histological response remains a strong prognostic factor for the cure fraction, indicating that patients who respond well to initial treatment are more likely to be cured. However, even after accounting for missing values, there is no clear indication that a good histological response is associated with longer progression-free survival (PFS) for uncured patients. This underscores the importance of having the possibility to differentiate between factors that influence the probability of cure from those that affect survival among uncured patients. Overall, we observe that, despite slight changes in the estimated coefficients and of the length of the confidence intervals, the conclusions about the significance of the covariate effects are the same for all the models and estimation procedures considered.
@@
As a further development of this work, one might consider an alternative way of performing multiple imputation based on the observed likelihood (instead of the complete one), which avoids considering the latent cure status as an additional variable with missing values. When there is only one missing variable, this would avoid the chained equations, since there would be no need to impute G. However, such an approach would require the estimation of the parameters $\alpha$ and $\beta$ in Step 2 of Algorithm 1 by fitting a mixture cure model and the variance needs to be estimated via computationally intensive bootstrap. Hence, we will explore this possibility in future works.
@@ B.2 Incomplete covariate only in the latency
We test the sufficient follow-up assumption for the osteosarcoma data analyzed in Section 5. Currently there exist no test uniformly over all covariate values so we apply the tests proposed by Xie et al. [2023] and Yuen and Musta [2024] separately for patients with good and poor histologic response, since that is the main prognostic factor for the cure probability.
The test proposed by Xie et al. [2023] considers the null hypothesis of sufficient follow-up $H_{0}:\tau_{F_{u}}\leq\tau_{C}$, where $\tau_{F_{u}}$ and $\tau_{C}$ denote, respectively, the right end points of the distribution of the event times for the uncured patients and of the censoring times. For patients with good histologic response, such test does not reject the null hypothesis of sufficient follow-up since $2(Y_{(n)}-\tilde{Y}_{(m)})<Y_{(n)}$, where $Y_{(n)}$ denotes the largest observed time and $\tilde{Y}_{(m)}$ denotes the largest observed event time. For patients with poor histologic response, the p-value of the test is 0.038 meaning that the null hypothesis is rejected at level 0.05.
However, it is argued by Yuen and Musta [2024] that one could instead test a more relaxed notion of practically sufficient follow-up. The null hypothesis is that the follow-up is insufficient $\tilde{H}_{0}:\tau_{C}\leq q_{1-\epsilon}$, where $q_{1-\epsilon}$ denotes the $(1-\epsilon)$-quantile of $F_{u}$ and $\epsilon$ is a small number chosen by the user. It essentially tests whether there is less that $\epsilon$ probability for the event to happen after the end of the study. In practice this would be considered as a sufficiently long follow-up if $\epsilon$ is around $1$-$2\%$. The test requires also the choice of a parameter $\tau$ such that the chances of the event happening after $\tau$ are negligible, i.e. a study of length $\tau$ would have been certainly sufficient. We apply the test for two different choices of $\epsilon$ and $\tau$ and p-values are given in Table 5. The results indicate that the follow-up is clearly sufficient for patients with good histologic response and borderline sufficient for those with poor histologic response.